OpenStudy (anonymous):

solve for y to 4 decimal places in; log base y for 10+log base 10 for y = 5

OpenStudy (anonymous):

\[\log_y(10)+\log_{10}(y)=5\] ?

OpenStudy (anonymous):

let's try this. by the change of base formula \[\log_y(10)=\frac{\log_{10}(10)}{\log_{10}(y)}=\frac{1}{log_{10}(y)}\]\]

OpenStudy (anonymous):

giving \[\frac{1}{\log(y)}+\log(y)=5\] and if we put \[\log(y)=x\] we have \[\frac{1}{x}+x=5\] \[1+x^2=5x\] \[x^2-5x+1=0\] \[x=\frac{5\pm\sqrt{21}}{2}\]

OpenStudy (anonymous):

now \[\log(y) = \frac{5+\sqrt{21}}{2}\] \[y=10^{\frac{5+\sqrt{21}}{2}}\] whatever that is. guess you need a calculator for that